OpenStudy (anonymous):

3(5/12)/1(2/3)=41/20 how?

OpenStudy (amistre64):

3(5/12) ------ = 3 * (15/24) if i see it right 1 (2/3)

OpenStudy (amistre64):

15/12 15/12 ----- = ------- = 15/8 2/3 * 4/4 8/12

OpenStudy (anonymous):

no i meant to say 3 and 5/12 being a fraction

OpenStudy (amistre64):

ohh ....

OpenStudy (amistre64):

3 + 5/12 = (36+5)/12 = 41/12

OpenStudy (amistre64):

1+2/3 = (3+2)/3 = 5/3

OpenStudy (amistre64):

41/12 41/12 ----- = ------ = 41/20 5/3 * 4/4 20/12

OpenStudy (anonymous):

(3*12+5)/12/(1*3+2)/3 =36+5/12/5/3 =41*3/12*5 =41/20

OpenStudy (anonymous):

atjari i dont see how you got that?

OpenStudy (phi):

can you change \[ 3 \frac{5}{12} \] into an improper fraction? you get \[ \frac{3*12+5}{12} = \frac{41}{12} \] now change the denominator into an improper fraction \[ 1 \frac{2}{3}= \frac{5}{3} \] to divide 41/12 by 5/3 , flip the 5/3 and multiply (important to know this!) \[\frac{41}{12}\cdot\frac{3}{5}= \frac{41}{\cancel{12}4}\cdot\frac{\cancel{3}}{5}= \frac{41}{20}\]

OpenStudy (anonymous):

Phi is right. That's hw I got t.

OpenStudy (anonymous):

yes thank yall so much